OpenStudy (anonymous):

Three waiters at a local restaurant earned different amounts of money the past week. Jennifer earned $40 less than twice the amount Tony earned. Chris earned $60 more than Tony. How much did each waiter earn if the total of all three was $900?

OpenStudy (anonymous):

\[(2T-40)+(T+60)+T= 900 \]\[4T+20=900\]subtract 20 on each side\[4T=880\]\[\frac{ 4T=880 }{ 4 }\]\[T=220\] Then you solve, tony makes $220 jennifer makes $400 chris makes $280